# Making Differential Privacy Easier to Use for Data Controllers and Data Analysts using a Privacy Risk Indicator and an Escrow-Based Platform

### v1

@@ Abstract
To address these two challenges: 1) we define a privacy risk indicator that indicates the impact of choosing $\epsilon$ on individuals’ privacy and use that to design an algorithm that chooses $\epsilon$ automatically; 2) we introduce a utility signaling protocol that helps analysts interpret the impact of DP on their downstream tasks. We implement the algorithm and the protocol inside a new platform built on top of a data escrow, which allows the controller to control the data flow and achieve trustworthiness while maintaining high performance. We demonstrate our contributions through an IRB-approved user study, extensive experimental evaluations, and comparison with other DP platforms. All in all, our work contributes to making DP easier to use by lowering adoption barriers.
1 Introduction
@@
Analyzing private data is essential for various tasks, and differential privacy (DP) is a promising tool to achieve privacy. In a typical DP deployment, data analyst runs queries on a dataset hosted on a platform that is accessible to a data controller. The dataset contains information about individuals, and releasing this information risks revealing the identity of those individuals. The data controller wants to protect their identity while releasing useful output to the data analyst. DP guarantees that the output of the query is likely to be the same if we change or remove any individual’s record in the dataset. This is achieved by adding calibrated noise to the output according to a privacy loss budget $\epsilon$: smaller $\epsilon$ leads to more accurate output but weaker privacy guarantees. Despite bespoke deployments by large organizations such as the US Census Bureau[3], Apple[5], and Google[24], it remains difficult to use DP in practice [23][43]. In this paper, we consider two important challenges controllers and analysts face.
@@
Challenge 1. The controller’s perspective. Controllers need to decide the level of privacy protection through $\epsilon$; however, it is challenging for the controller to reason about how much privacy each individual represented in the dataset receives with respect to $\epsilon$. Dwork et al.[23] showed that controllers have no clear consensus on how to choose $\epsilon$ or how to approach this problem in a principled manner.
@@
Our contribution. In this paper, we observe the crux of the problem is that $\epsilon$ does not directly translate into the degree of privacy loss incurred to specific individuals in the dataset. In response, we introduce a method to translate the effect of choosing $\epsilon$ into a privacy risk indicator (PRI) associated with each individual. Comparing any two individuals, the one with the higher PRI has greater privacy loss; thus, controllers learn how releasing a query output affects individuals’ privacy. And because the PRI and $\epsilon$ are connected in a principled way, there is an opportunity to automatically find an $\epsilon$ that produces a profile of privacy risk indicators consistent with the controller’s privacy requirement. We exploit this opportunity by designing an algorithm that automatically finds an $\epsilon$ and releases query output such that all individuals receive adequate protection based on the controller’s privacy requirement.
@@
Challenge 2. The analyst’s perspective. Analysts submit queries and receive noisy output that deviates from the ground truth output, i.e., the output if DP had not been applied. Without access to such ground truth, analysts do not know the impact of DP on the quality of their downstream tasks; this hinders the analysts’ ability to use the DP output in practice.
@@
Our contribution. In this paper, we propose a utility signaling protocol to help analysts interpret the impact of DP on their downstream tasks while still allowing controllers to fully control the data flow. In the protocol we propose, the analyst delegates their downstream task with an associated evaluation function to the platform. With the controller’s permission, the platform invokes the evaluation function on both the DP output and the ground truth output and releases both results, thereby signaling the utility loss on the analyst’s downstream task due to noise introduced by DP.
Putting everything together: Escrow-Based Platform. The utility signaling protocol requires analysts to place more trust in the platform than in a traditional DP deployment. We want to make the platform trustworthy while maintaining high performance, which is at odds with the increased overhead introduced by the algorithm we propose to help controllers. Our solution is to build a DP platform on top of a data escrow, which permits controlling data flow and improving trust with significantly lower overhead compared to other DP platforms.
Summary of related work. Many approaches have been proposed to help controllers choose $\epsilon$ [30] [26] [23] [36] [39] [40] [28] [48] [32] [28], [26] [30] [36] [39] [40], however, none concentrates on communicating the effects of $\epsilon$ on specific individuals represented in the dataset, as we do. In addition, approaches for communicating the privacy guarantees offered by DP are mainly targeted toward end-users[41][25], i.e., the privacy risk when sharing their data, whereas our approach focuses on helping controllers, who can only choose $\epsilon$ if they understand its privacy implications on those end-users.
Exposing error statistics, e.g., error bound, can help analysts interpret the output; such is the approach followed by the US Census [13]. However, such statistics are not sufficient in many cases. Our utility signaling protocol also complements existing approaches [47] to help analysts gain a deeper understanding of the effect of DP on their output.
Lastly, different deployment of DP has received a lot of attention [20] [42] [9] [4] [45] [46] [49] [10] [44] [11]. Our escrow-based implementation hits a new point in the design space between the degree of trust required by participants and the resulting performance.
@@
The rest of the paper is organized as follows. We present the new privacy risk indicator in Section 2 and the algorithm we designed to automatically find $\epsilon$ and release DP output in Section 3. Section 4 presents the utility signaling protocol and the escrow-based platform. We finish with evaluation (Section 5), related work (Section 6) and conclusions in Section 7.
2 Privacy Risk Indicator (PRI)
In this section, we present the new privacy risk indicator. We first provide background on differential privacy.
Differential privacy (DP) [19][21] is a mathematical definition of privacy that bounds the effect a single individual can have on the output of a query. Formally, a randomized mechanism $\mathcal{M}$ satisfies $\epsilon$-differential privacy if for any pair of neighboring datasets $x$ and $x^{\prime}$ where $x$ and $x^{\prime}$ differ by a single record, and for all sets of possible output $O\subseteq Range(\mathcal{M})$:
@@ 2 Privacy Risk Indicator (PRI)
The parameter $\epsilon$ is a tunable parameter called privacy loss budget. A typical way of achieving differential privacy is by adding noise to the output. The amount of noise is controlled by $\epsilon$; smaller $\epsilon$ leads to larger noise and less accurate output. Therefore, $\epsilon$ is the key parameter for navigating the privacy-accuracy trade-off in differential privacy.
We present two important properties of differential privacy:
Post-processing. [22] DP is immune to post-processing, which means any computation on the output of a DP mechanism will not undermine the DP guarantees. Formally, let $\mathcal{M}(x)$ be an $\epsilon$-DP mechanism. For any arbitrary function $f$, $f(\mathcal{M}(x))$ is $\epsilon$-DP.
Sequential composition. [22] Sequentially combining multiple DP mechanisms on the same input dataset results in quantifiable degradation of privacy. Formally, given $k$ independent mechanisms $\mathcal{M}_{1},\ldots,\mathcal{M}_{k}$ that satisfy $\epsilon_{1},\ldots,\epsilon_{k}$-DP, then any arbitrary function $g(\mathcal{M}_{1},\ldots,\mathcal{M}_{k})$ is $\sum_{i=1}^{i=k}\epsilon_{i}$-DP.
2.1 Overview
A good $\epsilon$ provides adequate privacy protection to all individuals while making the output still useful to analysts. However, it is inherently difficult to reason about the privacy implications of choosing $\epsilon$ for an specific individual in the dataset. Because $\epsilon$ bounds the privacy loss for any individual in expectation, but its concrete effect on a specific individual’s privacy is unknown. This is the source of difficulty for controllers whose ultimate goal is to protect individuals in the dataset.
We observe that controllers can only comfortably choose $\epsilon$ if they understand how each individual’s privacy is affected by that choice. To this end, we define a privacy risk indicator (PRI), which indicates the amount of privacy loss of each individual relative to others. For instance, a controller can set an $\epsilon$, compute the corresponding query output, and observe the resulting PRI for each individual. If the PRIs are equal, then all individuals will be equally protected. If they look different, that is an indication that certain individuals are at a higher risk of being identified and a suggestion for the controller to try a different $\epsilon$. The direct connection between the new PRI and differential privacy that we explain next is key to both providing a principled understanding of the new indicator and designing strategies to find $\epsilon$, including an algorithm that finds $\epsilon$ automatically, which we will present in Section 3.
2.2 Privacy Risk Indicator: Formal Definition
@@
We denote an $\epsilon$-DP mechanism as $\mathcal{M}:\mathcal{X}^{n}\times\mathcal{Q}\rightarrow\mathcal{Y}$, where $\mathcal{Q}$ is the query space and $\mathcal{Y}$ is the output space of $\mathcal{M}$. We consider numeric query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$ that maps a dataset to $k$ real numbers. Given a query $q\in\mathcal{Q}$ and a query output $r\sim\mathcal{M}(x,q,\epsilon)$ of mechanism $\mathcal{M}$ which takes the dataset $x$, the query $q$ and the privacy loss budget $\epsilon$ as input, we define the privacy risk indicator of individual $i\in[n]$ as follows:
@@
where $x_{-i}$ is the neighboring dataset of $x$ with individual $i$’s record $x_{i}$ removed. The difference is in the L1 norm.
Computing PRI is $\epsilon$-DP. Since $r$ is the output of an $\epsilon$-differentially private mechanism, computing PRI also satisfies $\epsilon$-differential privacy by DP’s post-processing property.
Connection between PRI and privacy loss. The PRI has a direct connection to the concept of privacy loss as defined in differential privacy. We formally show that when the Laplace Mechanism is used to compute the query output $r$, the connection is simple: larger PRI corresponds to larger privacy loss and vice versa.
First, we introduce the definition of privacy loss in DP[22].
@@
Given two neighboring dataset $x,x^{\prime}\in\mathcal{X}^{n}$ such that $\lVert x-x^{\prime}\rVert_{1}\leq 1$, for an output $r\sim\mathcal{Y}$, the privacy loss incurred upon observing $r$ is defined as:
@@
Next, we introduce the definition of Laplace Mechanism[22], a classic $\epsilon$-differentially private mechanism that we use to compute the query output $r$.
@@
where $Z_{i}\sim Lap(\frac{GS_{q}}{\epsilon})$ are i.i.d random variables drawn from the Laplace distribution  [^1] The global sensitivity of q in the $\ell_{1}$ norm is defined as $GS_{q}=\max|q(x)-q(x^{\prime})|$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$..
Using the Laplace Mechanism, the relative privacy loss of a specific individual $i$ can be expressed as:
@@
Since $\lVert r-q(x_{-i})\rVert_{1}$ is the definition of $PRI_{i}$, there exists a clear connection between $PRI_{i}$ and individual $i$’s privacy loss: larger PRI corresponds to larger privacy loss. Similarly to PRI, the privacy loss for a specific individual is data and query-dependent; essentially, it is the privacy loss individual $i$ will suffer if they participate in the analysis and the corresponding query output $r$ is released, compared to if they do not participate. Note that since we are using the Laplace Mechanism, which is $\epsilon$-DP, the privacy loss will always be bounded by $\epsilon$.
We have shown that the PRI indicates the relative privacy loss of each individual if a given output is released using DP. A data controller can leverage the PRI for many purposes; one natural goal is to exploit the relationship between PRI and DP to choose an $\epsilon$ and the corresponding output to release. In the next section, we explain different approaches the controller can use to achieve this goal and propose an algorithm.
3 Deriving Epsilon based on PRI
In this section, we describe how to leverage the PRI we introduced to address Challenge 1. We explain the two general approaches of using the PRI to find $\epsilon$ in Section 3.1. We describe the key observations we made to determine whether the PRIs offer adequate protection and the detailed algorithm for choosing $\epsilon$ and releasing query output in Section 3.2. In Section 3.3, we consider multiple queries and discuss how to bound the total privacy loss. Finally, we discuss the algorithm complexities and extensions in Section 3.4.
3.1 Equal PRI means Equal Protection
The goal of DP is to protect each individual equally in expectation, such that no individual suffers additional privacy loss than others by participating in the analysis; the privacy loss is bounded by $\epsilon$. The PRI indicates the effect of $\epsilon$ on each individual, thus becoming an actionable indicator that aids in selecting a suitable value. We describe two approaches to help controllers choose $\epsilon$ that leads to exactly equal or statistically similar PRIs. We show that both are impractical, which motivates our new algorithm (Section 3.2).
Figure 1: Example of using DP to equalize the PRIs. The goal is to find the largest $\epsilon$ that can make the PRIs equal.
Can we make the PRIs equal? Figure 1 illustrates this using an example dataset with four records. If the output was released directly without adding noise (thus no DP), some individuals would be more at risk than others: those with higher PRI, as shown in the leftmost table. By decreasing $\epsilon$, more noise is added to the output, so the PRIs become more similar. Then, to protect all individuals equally, the controller should choose the $\epsilon$ that makes the PRIs equal. Continuing with the example of Figure 1, setting $\epsilon$ to $0.5$ leads to equal PRIs among the four individuals while $\epsilon=1$ does not. When $\epsilon=0.2$, since it produces larger noise, the PRIs will remain equal, but it also leads to less accurate output. Therefore, $\epsilon=0.5$ is the best choice among the three values as it achieves the best privacy-accuracy trade-off.
Although it helps as an illustration, making PRIs exactly equal across individuals is generally infeasible unless an excessive amount of noise is added to the query output, which is undesirable because this means the output will not be useful for the analyst. It is also possible that a non-zero $\epsilon$ does not exist for which all PRIs are equal, and the only solution is to add an infinite amount of noise and make the PRIs all infinity.
Making PRIs statistically similar? One strategy is to consider the PRIs similar enough if they are statistically similar. Concretely, we can take two biased samples of PRIs: the set of the highest PRIs and the lowest PRIs. Then, we test the null hypothesis $H0$: the two samples come from the same distribution, using the statistical hypothesis framework. We consider the PRIs to be similar enough if we fail to reject $H0$.
Since we choose to compare two biased samples, any hypothesis tests that are sensitive to differences in locations (e.g., Mann-Whitney[37], Kolmogorov-Smirnov[29] or Mood’s Median Test) do not work well in practice since the ranks or the medians of the two samples will likely be very different, thus the tests will always tend to reject the null hypothesis. On the other hand, if we change the strategy to choosing random samples of PRIs and compare them, it becomes too easy to conclude the samples are similar. Therefore, an alternative strategy is needed to determine whether the PRIs provide adequate protection.
3.2 The Find-$\epsilon$-from-PRI Algorithm
Two key observations we make lead to our strategy for determining whether the PRIs provide adequate protection.
O1: Compute ratio between two PRIs. The first observation is we should compute the ratio between the PRIs of two individuals to determine whether they receive similar protection, since the comparison depends on the magnitude of the PRI. Consider two cases: two individuals with PRI of 2 and 1, and another two individuals with PRI of 10 and 9. In both cases, the difference between the two PRIs is 1, but the latter is preferable for equal protection since the ratio is close to 1.
Figure 2: Left tail with similar larger PRIs (green line) is better than right tail with similar smaller PRIs (red line)
O2: Left tail is better than right tail. The second observation is that if we cannot make the PRIs similar to each other as adding too much noise will make the output useless, it is better only to have left tail while making the larger PRIs similar than have right tail; because right tail means some individuals will have much higher risk than the rest, as shown in Figure 2. Even though the green line shows more individuals have relatively larger PRIs, their PRIs are similar enough so that they still receive similar protection; that is, they become indistinguishable with similar privacy loss.
These two key observations lead to a simple strategy for determining whether the PRIs provide adequate privacy protection. Based on O1, we can compute the ratio of the maximum PRI over the minimum PRI. Intuitively, as the ratio approaches 1, the PRIs become more similar and closer to the blue line shown in Figure 2. However, in many cases, this requires adding too much noise. Therefore, O2 allows the controller to trade privacy for more accuracy: instead of choosing the minimum PRI, the controller can choose the PRI at a specific percentile $p$. We fix the maximum PRI as the numerator to ensure the riskiest individuals will always be protected. The smaller $p$ is, the larger the ratio; therefore, more noise must be added (smaller $\epsilon$) to bring the ratio closer to 1. Although $p$ is an additional parameter the controller needs to set, it is more interpretable than $\epsilon$, since the controller only needs to decide at which point all the larger PRIs should be similar. For example, if the controller chooses the median (the 50th percentile), a ratio close to 1 indicates no individual is at considerably higher risk than the median individual.
Algorithm. Since PRI and $\epsilon$ are connected in a principled way, we can find a suitable $\epsilon$ that is consistent with the controller’s privacy requirement by computing the PRIs for different $\epsilon$, and choose the $\epsilon$ that is associated with PRIs such that the ratio between the maximum PRI and the PRI at the $p$th percentile is close to 1. The main intuition is to test different values of $\epsilon$ in descending order, and find the largest suitable $\epsilon$ that satisfies the controller’s privacy requirement while preserving as much accuracy as possible (note Figure 1).
Input : Dataset $x=(x_{1},\cdots,x_{n})$; query $q$; DP mechanism $\mathcal{M}$; consumed privacy budget $\epsilon_{c}$; percentile $p$; threshold $\tau$

### v2

@@ Abstract.
To address these two challenges: 1) we define a privacy risk indicator that indicates the impact of choosing $\epsilon$ on individuals’ privacy and use that to design an algorithm to choose $\epsilon$ and release output based on controllers’ privacy preferences; 2) we introduce a utility signaling protocol that helps analysts interpret the impact of DP on their downstream tasks. We implement the algorithm and the protocol inside a new platform built on top of a data escrow, which allows controllers to control dataflows while maintaining high performance. We demonstrate our contributions through an IRB-approved user study, extensive experimental evaluations, and comparison with other DP platforms. All in all, our work contributes to making DP easier to use by lowering adoption barriers.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
1. Introduction
@@
Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006b) is a promising tool to analyze data about individuals, but it remains challenging to use in practice. In a typical DP deployment, a data analyst runs queries on a dataset hosted on a platform that is accessible to a data controller. The dataset contains information about individuals, and releasing this information risks revealing the identity of those individuals. The data controller wants to protect individuals’ privacy while releasing useful output to the data analyst. DP guarantees that the output of the query is likely to be the same if we change or remove any individual’s record in the dataset. This is achieved by adding calibrated noise to the output according to a privacy loss budget $\epsilon$: smaller $\epsilon$ leads to more accurate output but weaker privacy guarantees. Despite some successful deployments of DP (Abowd, 2018; Apple, 2017; Erlingsson et al., 2014; Ding et al., 2017), a few roadblocks prevent broader DP deployment. In this paper, we consider three such roadblocks:
@@
Challenge 1: Choosing $\epsilon$. Controllers need to decide the level of privacy protection through $\epsilon$; however, it is challenging for the controller to reason about how much privacy each individual represented in the dataset receives with respect to $\epsilon$.  Dwork et al. showed that controllers have no clear consensus on how to choose $\epsilon$ or how to approach this problem in a principled manner.
@@
Our contribution. In this paper, we observe the crux of the problem is that $\epsilon$ does not directly translate into the degree of privacy loss incurred to specific individuals in the dataset. In response, we introduce a method to translate the effect of choosing $\epsilon$ into a privacy risk indicator (PRI) associated with each individual. Comparing any two individuals, the one with the higher PRI has greater privacy loss; thus, controllers learn how releasing a specific query output affects individuals’ privacy. And because PRI is connected to $\epsilon$, there is an opportunity to automatically find an $\epsilon$ that produces a profile of privacy risk indicators consistent with the controller’s privacy preferences; we exploit this opportunity by designing an algorithm, called Find-$\epsilon$-from-PRI, to choose $\epsilon$ and release DP output given a simple description of controllers’ privacy preferences.
@@
Challenge 2: Interpreting utility of DP output. Analysts submit queries and receive noisy output that deviates from the raw output, i.e., the ground truth output if DP had not been applied. Without access to such ground truth, analysts do not know the impact of DP on the quality of their downstream tasks, which is the reason why they sought the DP output in the first place. This hinders the analysts’ ability to use the DP output in practice.
@@
Our contribution. In this paper, we propose a utility signaling protocol to help analysts interpret the impact of DP on their downstream tasks while still allowing controllers to fully control data release. In the protocol we propose, the analyst delegates their downstream task with an associated evaluation function to the platform. With the controller’s permission, the platform invokes the evaluation function on both the DP output and the raw output and releases both results, thereby signaling the utility loss on the analyst’s downstream task due to noise introduced by DP.
Challenge 3: Deploying our solutions in a DP platform. Deploying our solutions to the previous two challenges brings a new challenge on the platform design. Specifically, implementing the new Find-$\epsilon$-from-PRI algorithm (addressing Challenge 1) and utility signaling protocol (addressing Challenge 2) calls for a platform that meets the four requirements: (1) it allows controllers to control dataflows, (2) it implements necessary security measures to protect against unintended dataflows, (3) it facilitates controllers to inspect dataflows, and (4) it adds little overhead. Various DP platforms have been proposed, but none meet all four requirements. For instance, the central DP model assumes adversaries are incapable of corrupting the platform and forcing unintended dataflows, which is unrealistic in practice (Near et al., 2021); platforms leveraging cryptographic primitives (Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018) make fewer assumptions about the adversaries, but often have significant computation overhead.
Our contribution. We propose a new platform built on top of a data escrow (Xia et al., 2022) which enables the implementation of the Find-$\epsilon$-from-PRI algorithm and utility signaling protocol, meeting the four platform requirements we curated. We show that the escrow-based platform hits a new point in the design space, offering a better trade-off between security, usability, and performance.
@@
Summary of Contributions and Related Work
Helping controller choose $\epsilon$ using PRI. Many approaches have been proposed to help controllers choose $\epsilon$ (Hsu et al., 2014; Ge et al., 2019; Dwork et al., 2019; Lee and Clifton, 2011; Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021; Ligett et al., 2017), however, none concentrate on communicating the effects of $\epsilon$ on specific individuals represented in the dataset, as we do. In addition, approaches for communicating the privacy guarantees offered by DP are mainly targeted toward end-users who share their data (Nanayakkara et al., 2023; Franzen et al., 2022), whereas our approach focuses on helping controllers, who can only make an informed choice of $\epsilon$ if they understand its privacy implications on those end-users.
Helping analyst via utility signaling protocol. Adding DP-style noise to the output affects the quality and usability of the data for the analyst’s downstream tasks (Hotz et al., 2022). For instance, the adoption of DP by the US Census has raised major concerns regarding the reliability of the census statistics after noise infusion and its effect on research and policy-making (Ruggles et al., 2019; Santos-Lozada et al., 2020; Kenny et al., 2021; ABOWD, [n.d.]). Our utility signaling protocol takes a step forward in helping analysts interpret the effect of DP on the released output.
Escrow-based DP platform. Various DP platforms have been proposed (Dwork et al., 2006a; Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018; Nguyen et al., 2020; Bittau et al., 2017), but none meets the requirements for deploying our solutions to help controllers and analysts. Our escrow-based platform meets those requirements, enabling practical deployment of DP.
Outline. The rest of the paper is organized as follows. We present the new privacy risk indicator in Section 2 and the algorithm we designed to automatically find $\epsilon$ and release DP output in Section 3. Section 4 presents the utility signaling protocol and Section 5 presents the escrow-based platform. We finish with evaluation (Section 6), related work (Section 7) and conclusions in Section 8.
2. Privacy Risk Indicator (PRI)
Background. Differential privacy (DP) (Dwork, 2006; Dwork et al., 2006b) is a mathematical definition of privacy that bounds the effect a single individual has on the output of a query over a dataset of records, where each record represents an individual. Formally, a randomized mechanism $\mathcal{M}$ satisfies $\epsilon$-differential privacy if for any pair of neighboring datasets $x$ and $x^{\prime}$ where $x$ and $x^{\prime}$ differ by a single record, and for all sets of possible output $O\subseteq Range(\mathcal{M})$:
@@ 2. Privacy Risk Indicator (PRI)
One way of achieving differential privacy is by constructing a mechanism that adds noise to the query output. The amount of noise is controlled by a tunable privacy loss budget parameter, $\epsilon$; smaller $\epsilon$ leads to larger noise and less accurate output: the choice of $\epsilon$ determines the privacy-accuracy tradeoff (Near et al., 2021).
2.1. Why Choosing $\epsilon$ is Hard and PRI Overview
Controllers decide what query output to release through $\epsilon$, which impacts individuals’ privacy. The parameter $\epsilon$ bounds the maximum privacy loss for any individual, but it is challenging for controllers to reason about the actual privacy implications of releasing an output for specific individuals in the dataset after choosing $\epsilon$ (Dwork et al., 2019).
We hypothesize that controllers can only comfortably choose $\epsilon$ if they understand how each individual’s privacy is affected by that choice. To this end, we define a privacy risk indicator (PRI), which indicates the amount of privacy loss of each individual relative to other individuals in the dataset, thus providing a direct indication of the privacy impact of releasing a query output. We construct PRI so that it is directly connected with the concept of privacy loss, which we explain next. This connection is key to both providing a principled understanding of the new indicator and designing strategies to find a suitable $\epsilon$. One strategy includes an algorithm to find $\epsilon$ based on the controller’s privacy preferences expressed in terms of PRIs, which we will explain in Section 3.
2.2. Privacy Risk Indicator: Formal Definition
@@
We denote an $\epsilon$-DP mechanism as $\mathcal{M}:\mathcal{X}^{n}\times\mathcal{Q}\rightarrow\mathcal{Y}$, where $\mathcal{Q}$ is the query space and $\mathcal{Y}$ is the output space of $\mathcal{M}$. Specifically, we consider numeric query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$ that maps a dataset to $k$ real numbers. Given a query $q\in\mathcal{Q}$ and a query output $r\sim\mathcal{M}(x,q,\epsilon)$ of mechanism $\mathcal{M}$ which takes the dataset $x$, the query $q$ and the privacy loss budget $\epsilon$ as input, we define the privacy risk indicator of individual $i\in[n]$ as follows:
@@
where $x_{-i}$ is the neighboring dataset of $x$ without $i$’s record $x_{i}$.
Connection between PRI and privacy loss. Next, we show that when the Laplace Mechanism (Dwork et al., 2014) is used to compute the query output $r$, PRI becomes a relative indicator of individual privacy loss. First, we show the definition of privacy loss in DP (Dwork et al., 2014).
@@
Given two neighboring datasets $x,x^{\prime}\in\mathcal{X}^{n}$ such that $\lVert x-x^{\prime}\rVert_{1}\leq 1$, for an output $r\sim\mathcal{Y}$, the privacy loss incurred upon observing $r$ is defined as:
@@
Next, we introduce the definition of Laplace Mechanism (Dwork et al., 2014), an $\epsilon$-differentially private mechanism that we use to compute the query output $r$.
@@
where $Z_{i}$ are i.i.d random variables drawn from the Laplace distribution with probability density function $p_{b}(z)=\frac{1}{2b}\exp(-|z|/b)$ and the scale $b=\frac{GS_{q}}{\epsilon}$  [^1] The global sensitivity of q in the $\ell_{1}$ norm is defined as $GS_{q}=\max|q(x)-q(x^{\prime})|$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$..
Using the Laplace Mechanism, we express the relative privacy loss of individual $i$ as:
@@
Claim: For all individuals $i,j$ where $i\neq j$, if $PRI_{i}>PRI_{j}$, then
$\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-i},q,\epsilon)}^{(r)}>\mathcal{L}_{\mathcal{M_{L}}(x,q,\epsilon)\parallel\mathcal{M_{L}}(x_{-j},q,\epsilon)}^{(r)}.$
The proof is straightforward: if $PRI_{i}>PRI_{j}$, then $\lVert r-q(x_{-i})\rVert_{1}$ $\lVert r-q(x_{-j})\rVert_{1}$, thus the inequality holds. Note that since we are using the Laplace Mechanism, which is $\epsilon$-DP, the privacy loss of any individual will always be bounded by $\epsilon$.
We now make several remarks:
PRI gives controllers additional information to decide what to release. The PRI becomes an additional indicator associated with each individual in the dataset that permits controllers to directly observe the effect of releasing a query output. In this paper, we show how PRI helps controllers decide which output to release.
PRI is data-dependent and thus leaks information about individuals. Controllers already have access to individuals’ raw data, so when they gain access to the PRI, they do not learn anything extra about the individuals. But this means the PRIs must remain private to the controllers and never be shared with anyone else.
Controllers may give PRI different uses. Controllers may express their privacy preferences for individuals in the dataset in terms of their associated PRIs, and derive an $\epsilon$ that produces that set of PRIs. This means that eliciting controllers’ privacy preferences based on PRIs paves the way for choosing $\epsilon$. We exploit this insight in the next section to design an algorithm to find $\epsilon$. We also discuss other ways in which PRIs may be used in Section 3.5.
3. Deriving Epsilon based on the PRI
In this section, we describe an algorithm to choose $\epsilon$ given controllers’ privacy preferences in terms of PRIs, thus addressing Challenge 1 (Sections 3.1, 3.2). In Section 3.3, we discuss how to bound the total privacy loss in the presence of multiple queries. In Section 3.4, we provide the end-to-end DP guarantee for the algorithm. We discuss the algorithm’s complexity and extensions in Section 3.5.
3.1. Using PRI to Express Privacy Preferences
Traditionally, controllers need to choose $\epsilon$ to convey their privacy preferences. However, choosing $\epsilon$ is challenging because it is hard to understand the specific effect of $\epsilon$ on the privacy of any one individual in the dataset. In contrast, the per-individual PRI indicates the relative privacy loss of an individual compared to other individuals in the dataset. The consequence is the PRI becomes an actionable indicator that helps controllers decide whether a concrete query output matches their privacy preferences, in which case they may agree to release the output. Next, we convey the intuition on how the PRI is useful to controllers through a vignette.
Figure 1. Example of using PRI to find $\epsilon$. For illustrative purposes, we show the PRI of each student (with the corresponding query output) and leave out students’ actual GPA.
Vignette illustrating one use of the PRI. Figure 1 shows the resulting PRIs of four students after answering a query about their average GPA. If the query output were released without using DP (a), the analyst observing the output would learn a lot about those students’ grades, as indicated by the PRI (remember that the PRI is a proxy to privacy loss). If the controller used DP, they would have to choose $\epsilon$ and obtain the query output by applying a DP mechanism. This would lead to a noisy output and a PRI associated with each student (note that running the mechanism two times with the same $\epsilon$ does not guarantee the same PRIs are obtained). We observe that decreasing $\epsilon$ (b, c, d) results in noisier output, which causes the PRIs to become more similar across students. In the extreme, when the PRIs are equal (c), no student in the dataset is at a higher risk than others. Thus, between (b) and (c), the controller may be more comfortable releasing the specific output (average GPA of 3.2) produced by choosing $\epsilon=1$ because it protects each student equally, which may be a particular privacy preference the controller wants to achieve. Lastly, the PRIs in both (c) and (d) indicate no student will be at a higher risk than others, but (d) produces a query output with worse accuracy. The controller may choose the larger $\epsilon$ (c) as it yields higher utility for the analyst receiving the output.
The above is an illustrative example: finding an $\epsilon$ that makes the PRI exactly equal across individuals is generally infeasible: either such an $\epsilon$ does not exist, or it does, but it requires adding too much noise that the output offers little utility to the analyst.
3.2. The Find-$\epsilon$-from-PRI Algorithm
When a controller runs a query on a dataset with a DP mechanism, observing the PRIs associated with individuals in the dataset communicates to the controller the relative privacy loss of each individual compared to others. Controllers could then change $\epsilon$, rerun the DP mechanism, and observe the resulting PRIs: they could repeat this process until they find a set of PRIs that represents their preferences. We introduce an algorithm that implements this intuition, named Find-$\epsilon$-from-PRI. In Section 3.1, we have shown one criterion for choosing $\epsilon$ that leads to exactly equal PRIs, which is typically infeasible. Instead, we propose an alternative criterion to let controllers decide whether a query output respects their privacy preferences. We explain the key idea and then the algorithm.
Key Idea. We let controllers express their privacy preferences by indicating a cutoff for the highest privacy loss of any individual in the dataset. This is done by requesting exactly one parameter from the controllers, a percentage p ($0\leq p\leq 100$). We compute the minimum and maximum PRI, $PRI_{min}$ and $PRI_{max}$. Then we find an $\epsilon$ that produces a list of PRIs (and associated DP output) such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$. In other words, we let controllers specify that the maximum PRI should deviate no more than $p$ percent from the minimum PRI. For instance, setting $p=0$ means the PRIs should be exactly equal.
Unlike $\epsilon$, the parameter $p$ directly conveys the level of risk conferred to individuals after releasing a specific query output. Smaller $p$ means more privacy at the cost of worse accuracy, hence controllers can trade privacy for accuracy by setting a larger $p$. The algorithm aims to find the largest $\epsilon$ that produces a list of PRIs such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$.
Input : Dataset $x=(x_{1},\cdots,x_{n})$; query $q$; percentage $p$